OpenStudy (anonymous):

Determine the solution set of ax 2 + bx + c = 0.

OpenStudy (anonymous):

Use the quadratic delta formula?

OpenStudy (akashdeepdeb):

x = [-b+- under root(b^2/4ac)]/2a :D

OpenStudy (anonymous):

can someone just solve it

OpenStudy (anonymous):

\[\Delta=b^2-4ac\] \[x_1=\frac{ -b-\sqrt{\Delta} }{ 2a }\] \[x_2=\frac{ -b+\sqrt{\Delta} }{ 2a }\]

OpenStudy (anonymous):

What is the solution set of 4x 2 - 11x + 6 = 0?

OpenStudy (anonymous):

It is {0.75, 2}

OpenStudy (anonymous):

Just learn the formula, and practice it, it is really easy.